OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick a green marble?

OpenStudy (anonymous):

\[\frac{3\text{ yellows}}{10\text{ total}}\cdot\frac{4\text{ greens}}{9\text{ total (no replacement)}}\]